OpenStudy (anonymous):

Identify all of the following solutions of square root of x minus 8 end root plus 8 equals x. x = 8 x = 9 x = 8 and x = 9 None of the above

OpenStudy (dumbcow):

\[x = \sqrt{x-8}+8\]

OpenStudy (aum):

\(\Large \sqrt{x-8} + 8 = x\) Quickest way would be to try out each answer choice,

OpenStudy (dumbcow):

old fashioned way -move 8 over -square both sides -solve quadratic

OpenStudy (anonymous):

so what is the answer?

OpenStudy (mathstudent55):

Above you were told how to solve this equation. What have you done so far?

OpenStudy (mathstudent55):

Of course, you can just substitute each choice in the equation and see which ones work.

OpenStudy (anonymous):

i would think its none of the above but idk

OpenStudy (mathstudent55):

Try them one by one. What do you get for the first one, x = 8?

OpenStudy (anonymous):

oh its 8 right?

OpenStudy (mathstudent55):